# Concept Steerers: Leveraging K-Sparse Autoencoders for Controllable Generations

### v1

@@ 4.3 Robustness to adversarial prompt manipulation
Table 3: Attack Success Rate (ASR) of different methods on various adversarial attack datasets. Lower ASR indicates better performance. Our method achieves the best overall robustness on average across all datasets by effectively removing nudity implicitly embedded in the model. Bold: best. Underline: second-best. Gray : require training and weight updates, Pink : do not require training but update model weights, Blue : do not require either.
Method Ring-A-Bell $\downarrow$ MMA-Diffusion $\downarrow$ P4D $\downarrow$ UnLearnDiffAtk $\downarrow$ Avg $\downarrow$
K77 K38 K16 Avg
SDv1.4 85.26 87.37 93.68 88.10 95.70 98.70 69.70 87.05
SA (Heng & Soh, 2024) 63.15 56.84 56.84 58.94 47.68 12.68 2.81 30.53
CA (Kumari et al., 2023) 86.32 91.69 94.26 90.76 10.60 5.63 1.04 27.01
ESD (Gandikota et al., 2023) 20.00 29.47 35.79 28.42 9.27 15.49 2.87 14.51
MACE (Lu et al., 2024) 2.10 0.00 0.00 0.70 2.72 2.82 1.51 1.94
UCE (Gandikota et al., 2024) 10.52 9.47 12.61 10.87 29.93 9.86 0.87 12.38
RECE (Gong et al., 2025) 5.26 4.21 5.26 4.91 21.77 5.63 0.72 8.76
SLD-Max (Schramowski et al., 2023) 23.16 32.63 42.11 32.63 35.76 9.14 2.44 20.24
SLD-Strong (Schramowski et al., 2023) 56.84 64.21 61.05 60.70 68.21 33.10 3.10 41.28
SLD-Medium (Schramowski et al., 2023) 92.63 88.42 91.05 90.70 68.21 24.00 1.98 46.72
SD-NP 17.89 40.42 34.74 31.68 24.00 10.00 1.46 16.29
SAFREE (Yoon et al., 2024) 35.78 47.36 55.78 46.31 40.82 10.56 1.45 24.29
TraSCE (Jain et al., 2024) 1.05 2.10 2.10 1.75 16.60 3.97 0.70 5.51
Ours 3.16 8.42 9.47 7.02 6.00 1.99 2.11 4.28
Next, we demonstrate the robustness of our method against adversarial prompts on four datasets: red-teaming approaches like Ring-A-Bell (Tsai et al., 2023), P4D (Chin et al., 2023), and attack frameworks like MMA-Diffusion (Yang et al., 2024) and UnlearnDiffAtk (Zhang et al., 2025). Adversarial prompts often consist of several non-English phrases or nonsensical text fragments that lack semantic meaning, but fool the underlying generative models to produce unsafe content. We follow the same setup in Sec. 4.1 and use a k-SAE trained on I2P prompts.
As shown in Table 3, our method achieves the best overall robustness on average across all datasets, significantly outperforming the most recent works TraSCE (Jain et al., 2024) by $1.23\%$ and SAFREE (Yoon et al., 2024) by 20.01%. Specifically, for the MMA-Diffusion and P4D datasets, our method achieves state-of-the-art results with improvements of $\mathbf{10.60\%}$ and $1.98\%$, respectively. This demonstrates that our method performs very well and can implicitly identify monosemantic interpretable directions for “nudity” within the latent space of adversarial prompts. Notably, our method outperforms RECE (Gong et al., 2025) specifically designed for tackling adversarial prompts by $\mathbf{4.48\%}$. For other datasets, our method ranks second-best or performs comparably to the best scores. We note that k-SAE is trained on text embeddings from I2P prompts to learn unsafe concepts and is not exposed to adversarial datasets. Remarkable performance in adversarial datasets demonstrates k-SAE generalizes well to unseen prompts, even without exposure to prompt embeddings from different distributions, similar observation to Sec. 4.2. We reiterate that once a k-SAE is trained on unsafe concepts, our method does not require retraining.
@@ 4.4 Efficiency of Concept Steerer
Table 4: Model Efficiency Comparison. Experiments were conducted on a single L40S GPU on P4D dataset (150 samples in total) for the task of removing nudity.
Method Inference Time (s/sample) $\downarrow$
SD 1.4 3.02
SAFREE (Yoon et al., 2024) 4.24
TraSCE (Jain et al., 2024) 15.62
@@
As shown in Table 4, our method achieves the fastest inference time among all other inference-based approaches, with only a $0.14$ sec./sample overhead on a single L40S GPU compared to the original SD 1.4. We highlight that our method is approximately $\mathbf{5}$x faster than the previous state-of-the-art (Jain et al., 2024) in unsafe concept removal.
4.5 Ablation Studies
Table 5: Attack Success Rate (ASR) when representations from different encoder layers are used to train k-SAE on the I2P dataset. The 10th layer yields the lowest ASR, indicating that this layer captures most information about nudity concept. k-SAE expansion factor = $4$, hidden neurons (n) = $3072$.
Layers ASR on I2P $\downarrow$
@@ 4.5 Ablation Studies
Table 6: Attack Success Rate (ASR) for different expansion factors of k-SAE trained on text embeddings extracted from the 10th layer of the I2P prompts. An expansion factor of 4 yields the lowest ASR, indicating its efficacy for steering.
[…]
@@
Table 7: Attack Success Rate (ASR) for different values of $\lambda$ of k-SAE with an expansion factor of $4$ trained on text embeddings of 10th layer on the I2P dataset. $\lambda=-0.5$ yields the lowest ASR.
$\lambda$ ASR on I2P $\downarrow$
$-0.1$ 2.59
$-0.2$ 1.23
$-0.3$ 0.87
$-0.4$ 0.60
$-0.5$ 0.36
Finally, we analyze the impact of our design choices on the overall steering capacity and visual quality.
Effect of Concept Steering on Visual Quality: To evaluate the impact of our approach on visual quality, we conduct a user study using $50$ randomly selected safe images generated by the original SD 1.4 model and nudity-steered images produced by applying our method on SD 1.4. We followed the setup described in Sec 4.1. The study involved $22$ participants, who were shown images in a randomized order and were asked to select the image they preferred most based purely on overall visual quality. $44.7\%$ of users preferred images produced by concept steering, while $44.9\%$ preferred images from SD 1.4, indicating that participants expressed an almost equal preference for both generations. This is a crucial finding because it shows that our method does not deteriorate visual quality from the base model but offers the additional benefit of controllability.
Effect of Layer Selection on Steering: We examine how the selection of different layers in the text encoder impacts the semantic information captured in k-SAE and thereby concept steering. As shown in Table 5, representations from later layers are more effective to remove nudity and steering than earlier layers. We believe that earlier layers capture more low-level semantic information, thus high-level concepts such as nudity are better captured in the later layers, making them suitable candidates for steering. Similar observations were reported in Toker et al. (2024).
Effect of k-SAE capacity on steering: We investigate the effect of k-SAE capacity determined by different expansion factors on steering results. From Table 6, we note that the performance differences between capacities is relatively minor, using an expansion factor of $4$ proves to be the most effective in removing nudity.
Effect of steering strength $\lambda$: Finally, we investigate the effect of the steering strength, $\lambda$. Table 7 illustrates the impact of $\lambda$, showing that decreasing its value enables more effective removal of nudity from a greater number of images. As shown in the first and second rows of Fig. 11, setting $\lambda=-0.1$ effectively removes the nudity concept in most images. However, smaller $\lambda$ values lead to a more complete removal, as demonstrated in the last row of Fig. 11.
@@ 5 Discussion and Future Work
We propose a novel framework leveraging k-SAEs to enable efficient and interpretable concept manipulation in diffusion models. Once trained, k-SAE serves as a Concept Steerer to precisely control specific visual concepts (e.g., nudity, violence, etc.) Our extensive experiments demonstrate that our approach is very simple, does not compromise the generation quality, and is robust to adversarial prompt manipulations. Currently, we steer concepts by extracting representations from the text encoder of the generative models. In future, we wish to explore steering via visual embeddings and allow users more control by selecting regions in an image and locally steer.
Impact Statement
As text-to-image models are increasingly integrated into high-stakes applications, discouraging unsafe generations is of paramount significance. This work presents an effective approach for identifying and suppressing unsafe concept directions across various generative models. By improving the controllability and reliability of generative models, our method advances the development of safer AI systems, facilitating their responsible deployment in real-world applications. Code is available at: https://github.com/kim-dahye/steerers
@@ Appendix A Implementation details
Training k-SAE with FLUX: For FLUX.1-dev (Labs, 2023) visualization, we train k-SAE using features extracted from the residual stream of the 23rd (out of 24) layer of the T5-XXL text encoder on prompts from the I2P dataset. The k-SAE is trained with $k=64$ and an expansion factor of $16$, resulting in a total hidden size dimension $n=65536$.
Text encoders of diffusion models: We extract text embeddings for k-SAE from CLIP ViT-L/14 (Radford et al., 2021) for SD 1.4, OpenCLIP-ViT/G (Cherti et al., 2023) and CLIP-ViT/L for SDXL-Turbo, and T5-XXL (Raffel et al., 2020) for FLUX.1-dev.
@@ Appendix B More details of the benchmarks
We evaluate our method for unsafe concept removal tasks on five publicly available inappropriate or adversarial prompts datasets following prior work (Jain et al., 2024): I2P [^3] https://huggingface.co/datasets/AIML-TUDA/i2p (Schramowski et al., 2023), Ring-A-Bell [^4] https://huggingface.co/datasets/Chia15/RingABell-Nudity (Tsai et al., 2023), P4D [^5] https://huggingface.co/datasets/joycenerd/p4d (Chin et al., 2023), MMA-Diffusion [^6] https://huggingface.co/datasets/YijunYang280/MMA-Diffusion-NSFW-adv-prompts-benchmark (Yang et al., 2024), and UnlearnDiffAtk [^7] https://github.com/OPTML-Group/Diffusion-MU-Attack/blob/main/prompts/nudity.csv (Zhang et al., 2025). I2P contains 4703 real user prompts that are likely to produce inappropriate images. Ring-A-Bell consists of two inappropriate categories: nudity and violence. For nudity, it contains 95 unsafe prompts for each split (K77, K38, and K16). For violence, we use the Ring-A-Bell Union dataset, which includes 750 prompts. P4D contains 151 unsafe prompts generated by white-box attacks on the ESD (Gandikota et al., 2023) and SLD (Schramowski et al., 2023). MMA-Diffusion contains 1000 strong adversarial prompts generated via a black-box attack. UnlearnDiffAtk contains 142 adversarial prompts generated using white-box adversarial attacks.

### v3

@@ 4.3 Robustness to adversarial prompt manipulation
Next, we demonstrate the robustness of our method against adversarial prompts on four datasets: red-teaming approaches like Ring-A-Bell tsai2023ring , P4D chin2023prompting4debugging , and attack frameworks like MMA-Diffusion yang2024mma and UnlearnDiffAtk zhang2025generate . Adversarial prompts often consist of several non-English phrases or nonsensical text fragments that lack semantic meaning, but fool the underlying generative models to produce unsafe content. We follow the same setup as in Sec. 4.1 and use a k-SAE trained on I2P prompts.
As shown in Table 1, our method achieves the best overall robustness on average across all datasets, significantly outperforming the most recent works TraSCE jain2024trasce by $1.23\%$ and SAFREE yoon2024safree by 20.01%. Specifically, for the MMA-Diffusion and P4D datasets, our method achieves state-of-the-art results with improvements of $\mathbf{10.60\%}$ and $1.98\%$, respectively. This demonstrates that our method performs very well and can implicitly identify monosemantic interpretable directions for “nudity” within the latent space of adversarial prompts. Notably, our method outperforms RECE gong2025reliable specifically designed for tackling adversarial prompts by $\mathbf{4.48\%}$. For other datasets, our method ranks second-best or performs comparably to the best ones. We note that k-SAE is trained on text embeddings from I2P prompts to learn unsafe concepts and is not exposed to adversarial datasets. Remarkable performance on adversarial datasets demonstrate that k-SAE generalizes well to unseen prompts, even without exposure to prompt embeddings from different distributions, as also observed in Sec. 4.2. We reiterate that once a k-SAE is trained on unsafe concepts, our method does not require retraining and can be applied at test-time.
@@ 4.4 Efficiency of Concept Steerer
Table 3: Model Efficiency Comparison. Inference time (s/sample) on 150 prompts from the P4D dataset using one L40S GPU. Lower is better.
Method Time $\downarrow$
SD 1.4 rombach2022high 3.02
SLD-Max schramowski2023safe 8.59
SAFREE yoon2024safree 4.24
TraSCE jain2024trasce 15.62
@@
As shown in Table 3, our method achieves the fastest inference time among all other inference-based approaches, with only a $0.14$ sec./sample overhead on a single L40S GPU compared to the original SD 1.4. We highlight that our method is approximately $\mathbf{5}$x faster than the previous state-of-the-art jain2024trasce in unsafe concept removal.
4.5 Analysis of Concept Steerer
Table 4: Attack Success Rate (ASR) on the I2P dataset. (a) ASR when representations from different encoder layers are used to train k-SAE. The 10th layer yields the lowest ASR, indicating that this layer captures most information about nudity concept. k-SAE expansion factor = 44, hidden neurons (n) = $3072$. (b) ASR for different expansion factors of k-SAE trained on text embeddings extracted from the 10th layer of the I2P prompts. An expansion factor of 4 yields the lowest ASR, indicating its efficacy for steering. (c) ASR for different values of $\lambda$ of k-SAE with an expansion factor of 4 trained on 10th-layer text embeddings. $\lambda=-0.5$ yields the lowest ASR.
Layer ASR $\downarrow$
@@ 4.5 Analysis of Concept Steerer
Expansion factor Capacity ASR $\downarrow$
@@
Below, we report the impact of different design choices of Concept Steerer, taking nudity removal as a sample task.
Effect of concept steering on visual quality: To evaluate this, we conduct a user study using $50$ randomly selected safe images generated by the original SD 1.4 model and nudity-steered images produced by applying our method on SD 1.4.  [^3] In our study, visual quality refers to how photo-realistic, visually clear, and pleasing an image is as perceived by the user. Interface screenshot in Appendix. We followed the setup described in Sec 4.1. The study involved $22$ voluntary participants, who were shown pairs of images in a randomized order and were asked to select the image they preferred most based purely on overall visual quality. $44.7\%$ of users preferred images produced by concept steering, while $44.9\%$ preferred images from SD 1.4, indicating that participants expressed an almost equal preference for both generations.44410.4% selected ”none of these” when they did not prefer either image. This is a crucial finding because it shows that our method does not deteriorate visual quality from the base model but offers the additional benefit of controllability.
We examine the effect of the choice of the text encoder’s layer on the semantic information captured in k-SAE and thereby concept steering. From Table 4(a), we note that representations from later layers are more effective at removing nudity (lower ASR) than earlier layers. We believe that earlier layers capture more low-level semantic information, thus high-level concepts such as nudity are better captured in the later layers, making them suitable candidates for steering. Similar observations were reported in toker2024diffusion .
We investigate the effect of k-SAE capacity determined by different expansion factors on steering results. From the ASR results in Table 4(b), we note that the performance differences between capacities is relatively minor, and an expansion factor of $4$ empirically is most effective in removing nudity.
We study the effect of steering strength $\lambda$ in Table 4(c), showing that decreasing its value enables more effective removal of nudity from a greater number of images. As shown in Fig. 8 (a), setting $\lambda=-0.1$ effectively removes the nudity in that example. However, smaller $\lambda$ values ($\lambda=-0.5$) lead to a more complete removal, as demonstrated in Fig. 8 (b) and lowest ASR scores in Table 4(c).
@@ 5 Discussion and Future Work
We propose a novel framework leveraging k-SAEs to enable efficient and interpretable concept manipulation in diffusion models. Once trained, k-SAE serves as a Concept Steerer to precisely control specific visual concepts (e.g., nudity, violence, camera angles etc.). Our extensive experiments demonstrate that our approach is very simple, does not compromise the visual quality, and is robust to adversarial prompt manipulations. Our current approach steers concepts by manipulating representations only from the text encoder. However, this design does not fully leverage the rich information encoded in the visual or the multi-modal embeddings of the diffusion model. Future work will explore interpreting and steering multi-modal embeddings to enable finer-grained control in generations. We note that sometimes higher steering strength values ($\lambda$) impact object and identity consistency, e.g., the style of the bench in Fig. 6. Exploring steering mechanisms where users select regions in an image to apply concept steering locally is a worthy direction to explore.
@@ Appendix A Implementation Details
Training k-SAE with FLUX: For FLUX.1-dev [59] visualization, we train k-SAE using features extracted from the residual stream of the 23rd (out of 24) layer of the T5-XXL text encoder on prompts from the I2P dataset. The k-SAE is trained with $k=64$ and an expansion factor of $16$, resulting in a total hidden size dimension $n=65536$.
Text encoders of diffusion models: We extract text embeddings for k-SAE from CLIP ViT-L/14 [60] for SD 1.4, OpenCLIP-ViT/G [61] and CLIP-ViT/L for SDXL-Turbo, and T5-XXL [62] for FLUX.1-dev.
Baselines: We compare our method against SDID [43], SLD [5], SD with negative prompt (SD-NP), SAFREE [9], TraSCE [10], ESD [6], FMN [15], CA [29], MACE [30], SA [7], UCE [41], and RECE [42]. We reported the performance of prior works with publicly available code. Whenever pretrained weights were provided, we directly used them, otherwise, we followed the publicly available implementations including hyperparameters as closely as possible. We also reported performance of several other methods as provided in the TraSCE [10], and we followed their experimental setup.
Experimental setup: When steering concepts, we use a slightly modified encoder of k-SAE without the TopK activation function, as the TopK function clamps some possible important information that is critical for both maintaining visual quality and effective concept removal. All experiments are conducted using Python 3.10.14 and PyTorch 2.5.1 [74], on a single NVIDIA L40 GPU with 48GB of memory.
@@ Appendix B More Details of the Benchmarks
We evaluate our method for unsafe concept removal tasks on five publicly available inappropriate or adversarial prompts datasets following prior work [10]. I2P [^5] https://huggingface.co/datasets/AIML-TUDA/i2p [5], Ring-A-Bell [^6] https://huggingface.co/datasets/Chia15/RingABell-Nudity [63], P4D [^7] https://huggingface.co/datasets/joycenerd/p4d [70], MMA-Diffusion [^8] https://huggingface.co/datasets/YijunYang280/MMA-Diffusion-NSFW-adv-prompts-benchmark [71], and UnlearnDiffAtk [^9] https://github.com/OPTML-Group/Diffusion-MU-Attack/blob/main/prompts/nudity.csv [72]. I2P contains 4703 real user prompts that are likely to produce inappropriate images. Ring-A-Bell consists of two inappropriate categories: nudity and violence. For nudity, it contains 95 unsafe prompts for each split (K77, K38, and K16). For violence, we use the Ring-A-Bell Union dataset, which includes 750 prompts. P4D contains 151 unsafe prompts generated by white-box attacks on the ESD [6] and SLD [5]. MMA-Diffusion contains 1000 strong adversarial prompts generated via a black-box attack. UnlearnDiffAtk contains 142 adversarial prompts generated using white-box adversarial attacks.
@@ Appendix C Additional Qualitative Results
Steering nudity concept on COCO dataset: To evaluate the effect of removing the nudity concept during the generation process, we apply different unsafe concept removal approaches to a safe dataset, i.e., COCO [69]. Figure 11 presents qualitative comparisons across methods. Our method generates images that are qualitatively similar to those from the original SD 1.4, even after removing the nudity concept, while preserving photo-realism and maintaining alignment with the input prompt. It achieves competitive results compared to other approaches, including SLD-Max [5], SAFREE [9], and TraSCE [10]. This highlights our method’s ability to selectively remove targeted concepts during generation without harming overall image quality or semantic fidelity.
@@
Appendix D Broader Impacts
As text-to-image models are increasingly integrated into high-stakes applications, discouraging unsafe generations is of paramount significance. This work presents an effective approach for identifying and suppressing unsafe concept directions across various generative models. By improving the controllability and reliability of generative models, our method advances the development of safer AI systems, facilitating their responsible deployment in real-world applications.
Appendix E User Study Interface
We provide the full instructions and a screenshot of the interface used in the user study described in Sec. 4.5. Participants were instructed as shown in Table 6, and the interface is illustrated in Fig. 17.
Appendix F Effect of Layer Selection on Steering
In Sec. 4.5, we studied the effect of layer selection on steering for the nudity concept. To examine whether this trend holds for a different concept, we analyze the effect of layer selection on steering for the violence concept. We report the Attack Success Rate (ASR) using representations from different encoder layers on the Ring-A-Bell-Union (Violence) dataset [63]. As shown in the Table 5, the 10th layer performs best (i.e., lowest ASR), although the difference with the 12th layer is marginal. Given the above results and those from Table 4 (a), we note that text representations of both violence and nudity concepts exhibit a similar trend, i.e., representations from later layers are more effective for steering compared to earlier layers. We attribute this to later layers encoding more complex, semantically rich information compared to earlier layers as also observed in  [73, 75, 76, 77].
Appendix G Impact of Different Steering Variants
In this section, we analyze the effect of alternative steering variants to elucidate the contribution of key components in our approach. Recall that our proposed formulation is $x_{\text{steered}}=x+W_{\text{dec}}(\lambda*\texttt{ENC}(x_{C}))$ as in Eq. 5 (L160-161). We stress that k-SAE is non-linear due to the use of ReLU and Top-K activation functions, as noted in Eq. 1 and L142-144.
To test the extent of non-linearity and its contribution to the learning of monosemantic concepts, we make an assumption that all components are linear. Under this assumption, $x_{\text{steered}}$ simplifies to $x+\lambda*W_{\text{dec}}(\texttt{ENC}(x_{C}))=x+\lambda*x_{C}+Error$, where $Error$ is the residual error needed to reconstruct $x_{C}$.
In fact, when $x_{C}$ is passed through the SAE, it can be approximated as $x_{C}\approx\texttt{DEC}(\texttt{ENC}(x_{C}))+Error,$ where $\texttt{DEC}(x)=W_{\text{dec}}x+b_{\text{pre}}$ as defined in Eq. 2. Substituting this in, we get $x_{C}\approx W_{\text{dec}}(\texttt{ENC}(x_{C}))+b_{\text{pre}}+Error$. Multiplying both sides by $\lambda$, we obtain $\lambda*x_{C}\approx\lambda*W_{\text{dec}}(\texttt{ENC}(x_{C}))+\lambda*b_{\text{pre}}+\lambda Error$. Under the linearity assumption, substituting back into the above equation gives: $x_{steered}=x+\lambda*W_{\text{dec}}(\texttt{ENC}(x_{C}))\approx x+\lambda*x_{C}-\lambda*b_{\text{pre}}-\lambda Error$.
To isolate the contribution of each component, we compare the following variants: (1) $x_{steered}=x+\lambda*x_{C},$ (2) $x_{steered}=x+\lambda*x_{C}-\lambda*b_{\text{pre}},$ and (3) $x_{steered}=x+\lambda*x_{C}-\lambda*Error$. These variants are derived from the approximate linear decomposition of the full formulation and allow us to evaluate the individual roles of directly injecting concept embedding $x_{C}$, pre-encoder bias term $\lambda*b_{\text{pre}}$, and residual error $\lambda*Error$, respectively. As shown in Fig. 18, these simplified variants fail to fully suppress the unsafe concept and often degrade visual quality, underscoring the effectiveness of our proposed steering formulation. This analysis reveals that simply injecting the concept embedding $x_{C}$, or removing the pre-encoder bias term or residual error, is insufficient for reliable concept erasure, and cannot be captured by a linear approximation, underscoring the necessity of our full formulation.
Appendix H License Information
We provide the license information for all assets used in this work. For additional details, please refer to the corresponding links.
Table 5: Attack Success Rate (ASR) when representations from different encoder layers are used to train k-SAE on the Ring-A-Bell-Union (Violence) dataset. The 10th layer yields the lowest ASR, indicating that this layer captures most information about nudity concept. k-SAE expansion factor = 44, hidden neurons (n) = $3072$.
Layers ASR $\downarrow$
12 43.86
10 43.73
8 83.86
6 82.13
4 85.46
@@ Appendix H License Information
Figure 11: Qualitative comparisons of different methods on the COCO dataset [69], including SLD-Max [5], SAFREE [9], and TraSCE [10]. After removing the nudity concept, our method produces images similar to SD 1.4 while better preserving prompt-image alignment and maintaining image quality. In contrast, other approaches often fail to preserve alignment and degrade image quality.
@@
Figure 15: Object attribute manipulation of SDXL-Turbo for the given prompt “A tree,” where the concept prompts are “An orange tree,” “A pink tree,” “A purple tree,” and “A yellow,” respectively. By adjusting $\lambda$, our method transitions the image toward the desired concept specified by the prompts.
Figure 16: Photographic style manipulation of SDXL-Turbo for the given prompt “A forest path,” where the concept prompts are “Foggy,” “Golden hour,” and “Sunset,” respectively. By adjusting $\lambda$, our method transitions the image toward the desired concept specified by the prompts.
You will see two images side by side. Your task is to choose the image you prefer (left or right) based on visual quality. In this study, ‘‘visual quality’’ refers to how realistic, clear, and visually pleasing an image appears.
Table 6: User study instruction.
Figure 17: User study interface. Our method (left) and SD 1.4 (right).
Figure 18: Effect of different steering methods on the P4D dataset. We compare three variants: (1) direct addition of the concept vector, (2) subtraction of the pre-encoder bias term, and (3) subtraction of the residual error. All variants are less effective than our proposed method, leading to either incomplete removal of unsafe content or degraded visual quality. These results highlight the importance of our approach.